OpenStudy (anonymous):

Find the maximum of g(x)=1+x+x2+x3+x4+x5 on the interval [0,1] .

OpenStudy (zarkon):

\(x,x^2,x^3,x^4,x^5\) are all increasing functions on [0,1] thus\( g(x)\) is an increasing function on [0,1] so it's min is at \(x=0\) and its max is at \(x=1\)

OpenStudy (anonymous):

It says that answer is incorrect for some reason.

OpenStudy (zarkon):

g(1)=6

OpenStudy (anonymous):

which equation did you use to get 6?

OpenStudy (zarkon):

\[g(x)=1+x+x^2+x^3+x^4+x^5\]

OpenStudy (anonymous):

oh i see! thank you so much! :)

OpenStudy (zarkon):

\[g(1)=1+1+1^2+1^3+1^4+1^5=1+1+1+1+1+1=6\]

OpenStudy (zarkon):

yw